# Safe and Optimal N-Spacecraft Swarm Reconfiguration in Non-Keplerian Cislunar Orbits

### v1

@@ Abstract
This paper presents a novel fuel-optimal guidance and control methodology for spacecraft swarm reconfiguration in Restricted Multi-Body Problems (RMBPs) with a guarantee of passive safety, maintaining miss distance even under abrupt loss of control authority. A new set of constraints exploits a quasi-periodic structure of RMBPs to guarantee passive safety. Particularly, this can be expressed as simple geometric constraints by solving optimal control in Local Toroidal Coordinates, which is based on a local eigenspace of a quasi-periodic motion around the corresponding periodic orbit. The proposed formulation enables a significant simplification of problem structure, which is highly applicable to large-scale swarm reconfiguration in cislunar orbits. The method is demonstrated in various models of RMBPs (Elliptical Restricted Three-Body Problem and Bi-Circular Restricted Four-Body Problem) and also validated in the full-ephemeris dynamics. By extending and generalizing well-known concepts from the two- to the three- and four-body problems, this paper lays the foundation for the practical control schemes of relative motion in cislunar space.
@@
$N$  = number of time discretization
@@
$\boldsymbol{r}_{ab}$  = position vector from $a$ to $b$
@@
$v/\mathcal{V}$  = TNW frame
@@ 1 Introduction
Distributed Space Systems (DSS) are rapidly emerging as a transformative paradigm in modern spaceflight, enabling new mission capabilities through spacecraft constellations, swarms, formations, and Rendezvous, Proximity Operations, and Docking (RPOD). Historically, almost all missions have primarily been executed in orbital regimes where the dynamics can be approximated as Keplerian with perturbations that are smaller than the gravitational force by magnitudes of order (e.g., Earth orbits, low lunar orbits). However, there is a surge of interest in deploying DSS in periodic orbits in the Restricted Multi-Body Problem (RMBP), where one such motivation comes from the projected deployment of NASA’s Gateway in a Near-Rectilinear Halo Orbit (NRHO) of the Earth-Moon $L_{2}$ libration point [2] and associated RPOD operations. Such a DSS requires Guidance, Navigation, and Control (GNC) algorithms that enable bounded relative motion between multiple spacecraft in the encountered multi-body dynamics.
@@
Traditionally, dynamics models and state representations for relative motion between spacecraft rely on a fundamental assumption of orbiting around a single primary (e.g., Earth) based on the Restricted Two-Body Problem (R2BP) with additional perturbations. These dynamics are often linearized around the chief orbit within Local-Vertical Local-Horizontal (LVLH) or Radial-Tangential-Normal (RTN) frames [3, 4]. In the last few decades, dynamics models in Relative Orbital Elements (ROE) [5] have also been developed as integration constants and fundamental matrix solutions of the aforementioned dynamics models, which has been demonstrated in numbers of flight missions, such as PRISMA [6], TanDEM-X [7], Starling [8], SWARM-EX [9], VISORS [10], and ADRAS-J [11]. Previous studies also include various perturbation sources such as geopotential spherical harmonics, solar radiation pressure, atmospheric drag, and third-body perturbations [12, 13]. Early research on relative motion with respect to a (virtual) spacecraft on a periodic orbit in the Restricted Three-Body Problem (R3BP) begins with the linearization of the nonlinear dynamics in the rotating frame [14, 15, 7], which naturally aligns with the formulation of the R3BP. More recently, relative dynamics of the Circular Restricted Three-Body Problem (CR3BP) and the Elliptic Restricted Three-Body Problem (ER3BP) have been derived in the LVLH frame [16, 17]. This is further extended to the Bi-Circular Restricted Four-Body Problem (BCR4BP) to express a higher-fidelity dynamics model in the Sun-Earth-Moon system [18].
@@
While some research introduces augmented dynamics with a continuous-time control law [15, 14, 19, 20] to realize a specific geometry of forced relative orbits, the majority of research in bounded relative motion in PRMBPs focuses on a fundamental solution space in the CR3BP. Besides five libration points and periodic orbit families surrounding them, equilibrium exists in the form of Quasi-Periodic Orbits (QPOs) that naturally trace the surface of a Quasi-Periodic Invariant Torus (QPIT) and envelop the nearby frequency-matching periodic orbit without an intersection [21]. Thus, a spacecraft on a QPO performs not only a quasi-periodic bounded motion to an attractor (e.g., Moon) in the rotating frame but also a quasi-periodic bounded relative motion with respect to a spacecraft in its co-moving orbital frame. For this reason, applications of QPOs to long-term formation-flight and proximity operations have been thoroughly investigated [22, 21, 17]. State-of-the-art demonstrations of QPO-based relative motion controls include a Linear Quadratic Regulator (LQR) controller and the Extended Kalman Filter for formation reconfiguration in the BCR4BP [23] or chance-constrained model predictive control (MPC) [24], relative station-keeping using invariant whiskers [25]. Recently, a new state representation for the relative motion with respect to a spacecraft on a CR3BP’s periodic orbit, known as Local Toroidal Coordinates (LTC) [22, 26], has been derived from the local eigenspace of a periodic orbit [27, 28]. The coordinates defined in this frame are fundamentally related to the geometric configuration of a torus, which offers a convenient first-order condition of the bounded relative motion between spacecraft in (quasi-)periodic orbits in the CR3BP. The LTC is applied to a long-term relative station-keeping of multi-spacecraft swarming in the CR3BP, which is then applied to the full-ephemeris dynamics of the Sun-Earth system [26]. However, formation reconfiguration in cislunar orbits becomes more complicated because of the larger dynamics modeling error (e.g., eccentricity of Earth-Moon orbit, Sun’s perturbations, etc.), as well as the fundamental complexity of the constrained trajectory planning compared to the unconstrained formation-keeping.
Rigorous safety criteria are one of the most critical considerations for bounded relative motion. Particularly, passive safety, which guarantees sufficient miss distance between spacecraft, even for free-drift trajectories that emanate from a sudden loss of control authority, is critical for practical operations in a chaotic dynamical system. Although the discussion of passive safety for relative motion in Perturbed Restricted Two-Body Problem (PR2BP) has been extensively studied in the literature [5, 29, 30, 31, 32], this has received little attention in the cislunar domain, with most discussions limited to topological analyses of the manifold structure [33, 34]. Two main issues need to be overcome in comparison to the PR2BP. The first is the choice of co-moving frame for resolving the relative dynamics in RMBPs. While the LVLH (RTN) frame is traditionally used [16], its tangential axis does not align with the chief’s velocity, especially in highly elliptical orbits such as NRHOs. This misalignment becomes critical when analyzing the spacecraft’s uncertainty ellipsoid, which is typically stretched along the velocity direction in the Earth-Moon rotating frame due to velocity-amplified positional uncertainty in navigation. As a result, the ellipsoid rotates within the LVLH frame throughout the orbit, complicating the geometric construction of safety regions such as Keep-Out Zones (KOZs). The second limitation is that the definition of passive safety is more convoluted in the RMBP. The difference in dynamical systems renders the direct application of the safety condition in the R2BP (such as relative eccentricity/inclination vector separation [35, 5, 36] in the ROE space) not attainable in PRMBPs. Also, there has been extensive works of passive safety that uses forward propagation of drift trajectories for a predefined period [29, 30, 31, 32, 37]. However, due to the high sensitivity and nonlinearity in PRMBPs, these methods require an extensive safety horizon, or the propagated set in a simplified dynamical model could be highly inaccurate in full-ephemeris dynamics. Therefore, deployment of the guidance and control strategy in not only the CR3BP but also in the other mid-fidelity dynamics models (e.g., ER3BP, BCR4BP) is of interest for the spacecraft swarm GNC design in the cislunar orbits.
@@
With these key limitations in mind, this paper proposes a passively safe reconfiguration of a spacecraft swarm in RMBPs, bridging the research gap between well-developed studies of bounded relative motion and control in the Perturbed Restricted Two-Body Problem (PR2BP) and those in RMBPs. Five main contributions are presented in this paper. First, the relative dynamics in the R3BP are newly derived in the TNW frame (“NTW frame" in Ref. 38, P.159), a co-moving frame defined based on the spacecraft velocity and angular momentum in the rotating frame. A QPO resolved in the TNW frame (defined as a Quasi-Periodic Relative Orbit (QPRO)) yields a compact bounded motion, containing large oscillatory behavior only in the velocity direction. This provides an easier understanding of relative motion in the RMBP compared to a description in the LVLH frame. Secondly, the geometric relationship between QPROs in the RMBP and periodic relative orbits in the R2BP is harmonized in this paper. It is shown that the local eigenspace associated with the periodic orbit in the Earth-Moon rotating frame defines the geometric configuration of the quasi-periodic relative motion. Moreover, insights into the minimum separation in relevant projections of the co-moving local frames are provided. Thirdly, the LTC is first established in an ER3BP’s sidereal-resonant periodic orbit and BCR4BP’s synodic-resonant periodic orbit, extended from the earlier work on relative motion in the CR3BP [22, 1] to a higher-fidelity dynamics model. Fourthly, an Optimal Control Problem (OCP) is solved in the LTC for the first time. Leveraging this state representation realizes a significantly simpler and more intuitive form of safety constraints. Furthermore, the proposed safety constraints are entirely decentralized for a given pair of relative orbits, which facilitates an extension to swarm control involving tens of spacecraft, while maintaining simplicity in onboard implementation. Finally, the optimized trajectory is deployed in the full-ephemeris model of the Sun-Earth-Moon system with an MPC scheme, providing a promising statistical result on the applicability of the proposed method in the full-ephemeris dynamics via Monte Carlo simulation. Overall, the proposed method enables the swarm reconfiguration in non-Keplerian cislunar orbits that is scalable to tens of spacecraft, which could open new possibilities of mission designs of cislunar RPOD or swarming.
@@
The remainder of the paper is organized as follows. First, nonlinear and linear equations of relative motion for the RMBP in the TNW frame are derived. Secondly, the passive safety of QRPOs is examined in relation to the relative orbit design in the PR2BP, followed by the derivation of the LTC. Third, using the invariance property of the local toroidal coordinates, an OCP is formulated as a nonconvex optimization problem and solved by Sequential Convex Programming (SCP). The control performances of the optimized trajectories are examined in the three dynamic models. Furthermore, the results of the OCPs are validated with full-ephemeris dynamics. Finally, conclusions and future research directions are provided.
@@ 1.1 Notations
2 Dynamics Models
This paper considers a spacecraft swarm comprising a (virtual) chief and $N_{sc}$ deputies, operating near the Moon and influenced by the gravitational forces of both the Earth and the Moon, and potentially by the Sun. In particular, RMBPs assume that the mass of each spacecraft is negligible compared to that of these celestial bodies [39]. The absolute positions of the chief and a deputy from the Moon’s center of mass are denoted as $\boldsymbol{r}$ and $\boldsymbol{r}_{d}$, respectively. The relative position of the deputy with respect to the chief is expressed as $\boldsymbol{\rho}=\boldsymbol{r}_{d}-\boldsymbol{r}$.
@@ 2.1 Coordinate Frames
The coordinate frames used in this article is summarized in Fig. 1. First, Fig. 1(a) illustrates absolute coordinate frames: inertial frame, barycenter synodic frame, and Moon synodic frame. The barycenter synodic frame is a rotating frame about the Earth-Moon barycenter $B_{1}$. The Moon synodic frame is attached to the barycenter synodic frame, while the origin of the frame is anchored to the Moon. In this paper, an absolute (reference) orbit/motion is defined as one that exhibits periodic motion around an attractor (e.g., Earth/Moon’s center of mass, libration points, etc.). These orbits/motions are usually resolved in the absolute frames.
Two co-moving coordinate frames are commonly used to describe the relative motion between spacecraft: LVLH and TNW frames, as shown in Figure 1(c). First, the LVLH frame $\mathcal{L}$ is defined by the right-handed triad $\{\hat{\boldsymbol{\imath}}_{L},\hat{\boldsymbol{\jmath}}_{L},\hat{\boldsymbol{k}}_{L}\}$ based on the absolute position of the chief spacecraft as $\hat{\boldsymbol{\imath}}_{L}=\hat{\boldsymbol{\jmath}}_{L}\times\hat{\boldsymbol{k}}_{L},\hat{\boldsymbol{\jmath}}_{L}=-{\boldsymbol{h}}/{h},\hat{\boldsymbol{k}}_{L}=-{\boldsymbol{r}}/{r}$. This triad can also define the Radial/Tangential/Normal (RTN) frame :$\mathcal{R}:\{-\hat{\boldsymbol{k}}_{L},\hat{\boldsymbol{\imath}}_{L},-\hat{\boldsymbol{\jmath}}_{L}\}$. The 2D planes defined using two of these axes are termed the RT, TN, and RN plane, respectively.
Similarly, the TNW frame $\mathcal{V}:\{\hat{\boldsymbol{\imath}},\hat{\boldsymbol{\jmath}},\hat{\boldsymbol{k}}\}$ is defined as a co-moving frame based on the spacecraft velocity as follows:
$\boldsymbol{\hat{\imath}}=\dfrac{\boldsymbol{v}}{v},\quad\boldsymbol{\hat{\jmath}}=\dfrac{\boldsymbol{h}}{h},\quad\boldsymbol{\hat{k}}=\boldsymbol{\hat{\imath}}\times\boldsymbol{\hat{\jmath}}.$ (1)
The 2D planes defined using two of these axes are termed TN, TW, and WN plane, respectively. The LVLH frame and the TNW frame coincide if an orbit is circular. Similarly to an absolute orbit, a relative orbit/motion is defined as a motion that exhibits (quasi-)periodicity around the chief spacecraft resolved in its co-moving frames.
(a) Absolute coordinate frames
@@
For the subscripts of an angular velocity $\boldsymbol{\omega}$ and a vector $[\boldsymbol{\zeta}]$, $b_{1}/\mathcal{B}$ denotes the barycenter synodic frame, $v/\mathcal{V}$ denotes the TNW frame, $m/\mathcal{M}$ denotes the Moon synodic frame, and $i/\mathcal{I}$ denotes the inertial frame.
2.2 Absolute Dynamics of RMBPs

### v2

@@ Abstract
This paper presents a novel fuel-optimal guidance and control methodology for spacecraft swarm reconfiguration in Restricted Multi-Body Problems (RMBPs) with a guarantee of passive safety, maintaining miss distance even under abrupt loss of control authority. A new set of constraints exploits a quasi-periodic structure of RMBPs to guarantee passive safety. Particularly, the condition for passive safety is expressed as simple geometric constraints by solving optimal control in Local Toroidal Coordinates, which is based on a local eigenspace of a quasi-periodic motion around the corresponding periodic orbit. The proposed formulation enables a significant simplification of problem structure, which is applicable to large-scale swarm reconfiguration in cislunar orbits. The method is demonstrated in the Circular Restricted Three-Body Problem, the Elliptic Restricted Three-Body Problem, and the Bi-Circular Restricted Four-Body Problem. Furthermore, the optimized control profiles are validated in the full-ephemeris dynamics model. By extending and generalizing well-known concepts of relative orbital elements within the restricted two-body problem to the three- and four-body problems, this paper lays the foundation for practical control schemes of relative motion in cislunar space.
@@
NN  = number of time steps
@@
$\boldsymbol{r}_{ab}$  = relative position vector from $a$ to $b$
@@
$v/\mathcal{V}$  = Velocity/Normal/Binormal (VNB) frame
@@ 1 Introduction
Distributed Space Systems (DSS) are rapidly emerging as a transformative paradigm in modern spaceflight, enabling new mission capabilities through spacecraft constellations, swarms, formations, and Rendezvous, Proximity Operations, and Docking (RPOD). To date, almost all RPOD missions have been executed in orbital regimes of the Perturbed Restricted Two-Body Problem (PR2BP), where Keplerian motion dominates and perturbations are several orders of magnitude weaker (e.g., Earth orbits, low lunar orbits). Recently, there has been growing interest in deploying DSS along periodic orbits governed by the Restricted Multi-Body Problem (RMBP). One such motivation comes from the projected deployment of NASA’s Gateway in a Near-Rectilinear Halo Orbit (NRHO) of the Earth-Moon $L_{2}$ libration point [2] and associated RPOD operations. Such a DSS requires Guidance, Navigation, and Control (GNC) algorithms that enable bounded relative motion between multiple spacecraft in the encountered multi-body dynamics.
@@
Traditionally, dynamics models and state representations for relative motion between spacecraft rely on a fundamental assumption of orbiting around a single primary (e.g., Earth) based on the PR2BP. These dynamics are often linearized around the chief orbit within Local-Vertical Local-Horizontal (LVLH) or Radial-Tangential-Normal (RTN) frames [3, 4]. In the last few decades, dynamics models in Relative Orbital Elements (ROE) [5, 6], which are based on the difference in the Keplerian orbital elements, have also been developed as integration constants and fundamental matrix solutions of the aforementioned dynamics models. The ROE-based GNC algorithms have been demonstrated in a number of flight missions, such as GRACE [7], PRISMA [8], TanDEM-X [9], AVANTI [10], Starling [11], SWARM-EX [12], VISORS [13], and ADRAS-J [14]. Previous studies also include various perturbation sources such as geopotential spherical harmonics, solar radiation pressure, atmospheric drag, and third-body perturbations [15, 16]. Early research on relative motion between spacecraft on a periodic orbit in the RMBPs begins with the linearization of the nonlinear dynamics in the rotating frame [17, 18, 9]. More recently, relative dynamics of the Circular Restricted Three-Body Problem (CR3BP) and the Elliptic Restricted Three-Body Problem (ER3BP) have been derived in the LVLH frame [19, 20]. This is further extended to the Bi-Circular Restricted Four-Body Problem (BCR4BP) to express a higher-fidelity dynamics model in the Sun-Earth-Moon system [21]. Furthermore, the relative dynamics of the CR3BP resolved in a co-moving frame based on spacecraft velocity (“NTW frame" in Ref. 22, “velocity-reference frame" in Refs. [23, 24], “Velocity/Normal/Conormal (VNC) frame" in Ref. [25], and “Velocity/Normal/Binormal (VNB) frame" in NASA General Mission Analysis Tool (GMAT) [26]) are derived in Refs. [1, 24] contemporaneously.
@@
Stable solutions in the CR3BP provide crucial insights into the design of bounded relative motion in PRMBPs. Besides five libration points and periodic orbit families surrounding them, equilibria exist in the form of Quasi-Periodic Orbits (QPOs) that naturally trace the surface of a Quasi-Periodic Invariant Torus (QPIT) and envelop the nearby frequency-matching periodic orbit without an intersection [27, 28, 29]. Thus, a spacecraft on a QPO performs not only a quasi-periodic bounded motion to an attractor (e.g., Moon) in the rotating frame but also a quasi-periodic bounded relative motion with respect to a spacecraft in its co-moving orbital frame. For this reason, applications of QPOs to long-term formation-flight and proximity operations in the RMBPs have been thoroughly investigated [27, 30, 31]. The characteristics of relative orbital motion are generalized by Lyapunov-Floquet theory [32] with analysis of the eigensystem of the dynamics. In the PR2BP, six unit eigenvalues of the monodromy matrix reflect the six first integrals (Integration Constants (ICs) [33, 34]) that lead to Relative Orbital Elements (ROEs) [5, 35, 36]. Meanwhile, in the PMBPs, the monodromy matrix and Floquet matrix are used to analyze the stability of the (quasi-)periodic orbits to analyze the fundamental structure of the quasi-periodic solutions [28, 29, 37]. Originally devised for different dynamical regimes, the two approaches converge within the methods of fundamental modal solutions that represent relative motion between spacecraft as a superposition of modes whose coefficients are prescribed by the ICs [38]. The ICs are derived for the relative motion in CR3BP, and the impulsive control scheme is solved within this state space [24]. Although this method provides a mathematically convenient structure for discussing the natural perturbation, a set of ICs (ROEs) lacks geometric insight into the spacecraft formation. Recently, a new state representation, Local Toroidal Coordinates (LTC) [39, 40], has been proposed to relate the dynamical motion to the geometry of the torus and offer convenient first-order conditions for bounded relative motion between spacecraft. Previous works adopted the LTC to perform a long-term relative station-keeping of multi-spacecraft swarming in the CR3BP, which is then applied to the full-ephemeris dynamics of the Sun-Earth system [40]. However, multi-revolution formation reconfiguration in cislunar orbits becomes more complicated due to a larger dynamics modeling error (e.g., eccentricity of Earth-Moon orbit, Sun’s perturbations, etc.), as well as the fundamental complexity of the constrained trajectory planning compared to the unconstrained formation-keeping.
Rigorous safety criteria are one of the most critical considerations for bounded relative motion. Particularly, passive safety, which guarantees sufficient miss distance between spacecraft, even for free-drift trajectories that emanate from a sudden loss of control authority, is critical for practical operations in a chaotic dynamical system. However, this has received little attention in the relative motion control in multi-body dynamics. Recent works discuss the safety of quasi-periodic relative orbits by computing the minimum separation between a QPO and the corresponding periodic orbit in the CR3BP using the LTC [39] as well as the fundamental modal solution space [24]. Particularly, Ref. [24] performs its analysis in the velocity-referenced frame, noting that the local-horizontal axis of the LVLH frame departs from the velocity direction on highly elliptic orbits such as NRHOs. Aligning the frame with the velocity direction keeps the navigation-error ellipsoid naturally elongated along its axis, which simplifies the construction of geometric safety regions of Keep-Out Zones (KOZs). Besides, formation-keeping control strategies that can target specific regions of the local eigenspace using both impulsive control [41, 31] and continuous control [18, 17, 42, 43] are developed to realize a specific geometry with some separations. In contrast, some works bypass the natural dynamics structure and perform reconfigurations between safe QPROs with a Linear Quadratic Regulator (LQR) controller in the BCR4BP [44] or Model Predictive Control (MPC) with Cartesian collision-avoidance constraints in the CR3BP [45, 46]. Nonetheless, the major limitation persists as a design that maintains passive safety throughout the transfer is more convoluted than safe orbit designs. Passive safety in the relative motion in the PR2BP has been mainly addressed by careful relative orbit design that ensures planar separation [7, 5], or forward/backward propagation of drift trajectories over a fixed duration [47, 48, 49, 50, 51]. In RMBPs, the former is difficult due to distinct dynamical structures. The latter approach has been applied to cislunar rendezvous [52], although this is also limited by the need for a long safety horizon due to nonlinear sensitivity in PRMBPs [52] and by the quadratic growth of constraints with the number of spacecraft. Consequently, there is a need for scalable safety-preserving constraints that extend beyond the CR3BP to mid- and high-fidelity relative dynamics models in RMBPs.
@@
With these key limitations in mind, this paper proposes a passively safe, fuel-optimal reconfiguration of a spacecraft swarm in RMBPs, bridging the research gap between well-developed studies of bounded relative motion and control in PR2BP and those in RMBPs. Five main contributions are presented in this paper. First, the relative dynamics in the RMBPs and the corresponding torus structures in the co-moving frame are compared between the LVLH frame and the velocity-reference frame. The analysis presents that the QPO resolved in the co-moving frame (termed as a Quasi-Periodic Relative Orbit (QPRO)) yields a compact bounded motion, containing large oscillatory behavior only in the velocity direction. This provides an easier understanding of relative motion in the RMBPs compared to a description in the LVLH frame. Secondly, the geometric relationship between QPROs in the RMBPs and periodic relative orbits in the PR2BP is harmonized in this paper. It is shown that the local eigenspace associated with the periodic orbit in the Earth-Moon rotating frame defines the geometric configuration of the quasi-periodic relative motion. Moreover, insights into the minimum separation in relevant projections of the co-moving local frames are provided. Thirdly, the LTC is constructed on a sidereal-resonant periodic orbit in the ER3BP and a synodic-resonant periodic orbit of the BCR4BP, extended from earlier works on relative motion in the CR3BP [39, 1] to a higher-fidelity dynamics model. Fourthly, an Optimal Control Problem (OCP) is solved in the LTC for the first time. Leveraging this state representation realizes a significantly simpler and more intuitive form of safety constraints. Furthermore, the proposed safety constraints are entirely decentralized for a given pair of relative orbits, which facilitates an extension to swarm control involving tens of spacecraft while maintaining simplicity in onboard implementation. Finally, the optimized controls are deployed in the full-ephemeris model of the Sun-Earth-Moon system with an MPC scheme, providing a promising statistical result on the applicability of the proposed method in the full-ephemeris dynamics via Monte Carlo simulation. Overall, the proposed method enables the swarm reconfiguration in non-Keplerian cislunar orbits that is scalable to tens of spacecraft, which could open new possibilities for mission designs of cislunar RPOD or swarming.
@@
The remainder of the paper is organized as follows. First, nonlinear and linear equations of relative motion for the RMBPs in the velocity-reference frame are derived. Secondly, the passive safety of QRPOs is examined in relation to the relative orbit design in the PR2BP, followed by the derivation of the LTC. Third, using the invariance property of the LTC, an OCP is formulated as a nonconvex optimization problem and solved via Sequential Convex Programming (SCP) [53]. The control performances of the optimized trajectories are examined in the three dynamic models. Furthermore, the results of the OCPs are validated with full-ephemeris dynamics. Finally, conclusions and future research directions are provided.
2 Dynamics Models
This paper considers a spacecraft swarm comprising a (virtual) chief and $N_{sc}$ deputies, operating near the Moon and subject to gravitational forces of multiple attractors, including the Earth, Moon, and Sun. In particular, RMBPs assume that the mass of each spacecraft is negligible compared to that of these celestial bodies.
@@ 2.2 Coordinate Frames
The coordinate frames used in this article is summarized in Fig. 1. First, Fig. 1(a) illustrates coordinate frames within the Earth-Moon system: inertial frame $\mathcal{I}$, barycenter synodic frame $\mathcal{B}$, and Moon synodic frame $\mathcal{M}$. The barycenter synodic frame is a rotating frame about the Earth-Moon barycenter $B_{1}$. The Moon synodic frame is attached to the barycenter synodic frame, while the origin of the frame is anchored to the Moon’s center of mass.
To describe the relative position and velocity between spacecraft near a (quasi-)periodic orbit (called reference orbit) in the synodic frame, two co-moving frames are employed: LVLH frame $\mathcal{L}=\{\hat{\imath}_{L},\hat{\jmath}_{L},\hat{k}_{L}\}$ and VNB frame $\mathcal{V}:\{\hat{\boldsymbol{\imath}},\hat{\boldsymbol{\jmath}},\hat{\boldsymbol{k}}\}$, as shown in Fig. 1(c). The LVLH frame is defined based on the position of the spacecraft as
$\hat{\boldsymbol{\imath}}_{L}=\hat{\boldsymbol{\jmath}}_{L}\times\hat{\boldsymbol{k}}_{L},\quad\hat{\boldsymbol{\jmath}}_{L}=-(\boldsymbol{r}\times\boldsymbol{v})/\|\boldsymbol{r}\times\boldsymbol{v}\|,\quad\hat{\boldsymbol{k}}_{L}=-{\boldsymbol{r}}/{r},$ (1)
whereas the VNB frame is defined based on the velocity of the spacecraft as
$\boldsymbol{\hat{\imath}}=\boldsymbol{v}/{v},\quad\boldsymbol{\hat{\jmath}}=(\boldsymbol{r}\times\boldsymbol{v})/\|\boldsymbol{r}\times\boldsymbol{v}\|,\quad\boldsymbol{\hat{k}}=\boldsymbol{\hat{\imath}}\times\boldsymbol{\hat{\jmath}}$. (2)
Here, $\boldsymbol{v}$ is the velocity vector resolved in the synodic frame. The LVLH axes are labeled as X, Y, and Z axes, forming the XY, YZ, and ZX planes. Similarly, the VNB frame uses the V, N, and B axes, which define the BV, VN, and NB planes.
(a) Synodic and inertial frames
@@
2.3 Dynamics of RMBPs